Seth transport operates a fleet of Motor Lorries. The record for Lorry-4150 reveals the following information Days Maintained 30 Days operated 25 Days Idle 5 Total hours operated 300 Total kms covered 2500 Total tonnage carried 200 (4 tonne – load per trip, return journey empty) Total cost for the month 270000 Prepare a performance statement reflecting a. Cost per day operated, cost per kilometer, cost per hour b. Cost per round trip, Tonne Km, cost per commercial tonne –km.


1
Expert's answer
2021-04-08T05:28:42-0400

a.


"\frac {270000}{30}=90000"

"\frac {270000}{2500}=108"

"\frac {270000}{300}=900"

b.


"\frac {2500}{200}=12.5"

"\frac {270000}{12.5}=21600"

"\frac {1250}{200}=6.25"

"\frac {270000}{6.25}=43200"